Operations Report – December 3rd

SOL: 6

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: Yes
Hours: 202.0 (Start), 203.2 (End)
Beginning charge: 100%
Ending charge: 53%
Currently charging: No

Opportunity rover used: Yes
Hours: 106.2 (Start), 107.3 (End)
Beginning charge: 100%
Ending charge: 50%
Currently charging: No

Curiosity rover used: Yes
Hours: 214.5 (Start), 215.3 (End)
Beginning charge: 100%
Ending charge: 54%
Currently charging: No

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 175 gallons remaining (375 gallons used)
Water (static tank): A “Supplemental Operations Report” was received by the crew today from Dr. Shannon Rupert. She informed us that our water tank is at 175 gallons, and that a water resupply vessel will arrive tomorrow to add water to the tank.
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes. The tank was emptied this morning at ~7:45 AM.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer and Commander inspected all six of the two-piece EVA suits today by opening the backpacks and checking the battery voltage with the multimeter. Three of the six suits were reading a voltage much lower that ~12 V despite being plugged in properly (switch was in the “off” position) and charging all night.

Thus, we believe that the three two-piece suits reading low voltages need new batteries. Please advise if you would like the Crew Engineer to locate and replace the batteries on EVA Suits #3, #4, and #6.

Furthermore, we decided to test out the performance of one of the functional suits to ensure that it could withstand a three-hour EVA. To that end, we turned on EVA Suit #2 at 3:28 PM and let it run for three hours. The suit remained functional for the full three-hour time block. Thus, we can confidently conclude that EVA Suits #1, #2, and #5, the suits which had > 12 V readings, are suitable for EVA use.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~31 gallons (~3 gallons today)
Heater: On
Supplemental light: In the “Supplemental Operations Report” we received today from Dr. Shannon Rupert, she said that the supplemental lighting will now be activated from 10 PM to 2 AM. We will inform Mission Support if we notice the supplemental light on at any other times.
Harvest: N/A

Summary of ScienceDome operations: No operations performed today by the crew in the Science Dome.
Dual split: In the “Supplemental Operations Report” we received today from Dr. Shannon Rupert, she said that she is turning on the heat in the Science Dome beginning today. We acknowledge the note to not adjust the settings in any way.

Summary of RAM operations: N/A

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: There is a lingering odor coming from the bathroom. The Crew Engineer, with the assistance of the GreenHab Officer, emptied the tank at ~7:45 AM this morning, yet the odor is still present ~11 hours later. The handbook mentions that we should consult with Mission Support if this is the case. Please advise us on how to improve this odor situation. We look forward to receiving your suggestions.

Copyright © The Mars Society. All rights reserved. | Main Site